OpenStudy (anonymous):

Help with this please. Find the product mentally. [(x + 1)(x - 1)]2 = 1. x4 + 2x2 - 1 2. x2 - 1 3. x4 - 2x2 + 1 Help please!

OpenStudy (anonymous):

@phi @dumbcow any help please?

OpenStudy (phi):

They want you to memorize how to factor a "difference of squares" Does that ring any bells? How do you factor \( (a^2 - b^2) \)?

OpenStudy (phi):

and they also want you to know how to expand the square \[ (a-b)2 \] which has a very simple pattern (that they want you to memorize... so you can just write it down)

OpenStudy (anonymous):

Well it would be... (a + b) (a - b) ?

OpenStudy (phi):

yes. Now look at your problem.... you are going the "other way"

OpenStudy (anonymous):

So instead it would be... (a - b) (a + c) or is it going to be.. (b + a) ( b - a) ?

OpenStudy (phi):

(x+1)(x-1) that is like (a+b)(a-b) what does it multiply out to ?

OpenStudy (anonymous):

x^2 + 1 x + 1x -1? Is that correct? Sorry I am horrible with math

OpenStudy (anonymous):

So then that would be "x^2+2x-1?"

OpenStudy (phi):

you know (a^2 - b^2) = (a+b)(a-b) that works both ways: (a+b)(a-b) = (a^2 - b^2)

OpenStudy (phi):

so you should look at (x+1)(x-1) and say "aha! those are the factors for a difference of squares"

OpenStudy (anonymous):

but wait it's squared as a whole! ahaaaa. So does that mean it would be x^4+2x^2-1....?

OpenStudy (phi):

the idea is in steps: (x+1)(x-1) is (x^2 - 1) but we want \( (x^2-1)^2 \) we use this pattern: \( (a-b)^2 = a^2 -2ab + b^2 \) can you see what the answer is (with a being x^2 and b being 1)